# Bridging the Social & Technical Divide in Augmentative and Alternative Communication (AAC) Applications for Autistic Adults

### v1

@@ 25-34 Semi-speaking 5-9 Proloquo4Text,
25-35 Mostly non-speaking $\leq$ 1 Proloquo,
@@ 5. Results
This section delves into the categories identified during the qualitative analysis of the transcripts. We will use a numerical value enclosed in parentheses is provided to indicate the number of participants supporting that particular theme (e.g., ”(3)” signifies ”three interviewees”). We will not be using any individual IDs due to the close-knit community of AAC users. We will also be referring to everyone by ”they/them”, even though we collected their pronouns during the interview. Direct quotes will be found inside quotation marks and italicized. The complete list of all 12 participants can be found in Table 1.
We have identified 9 main categories of themes, which we will describe in more detail in this section. The categories are: input options (§5.1), output options (§5.2), selecting or adapting AAC for a good fit (§5.3), when to start or swap AAC (§5.4), benefits (§5.5), access (§5.6), stumbling blocks for continued use (§5.7), social concerns (§5.8), and lack of control (§5.9).
5.1. Input Options
All 12 participants had preferences in terms of what input type they preferred or would like to see as an input option in a future AAC application. In this section, we discuss the use of keyboards, buttons/symbols, and other more indirect types of selection in applications, the differences felt between using symbol-based and text-based AAC, and aspects that make input easier such as vocabulary & layout flexibility, predictive text, or pre-programmed phrases.
@@ 5.1.1. Typing
All participants (12) wanted some type of typing input option to use at least occasionally, even if it is to augment their vocabulary for when they don’t have an appropriate button. Although a few (3) prefer a physical keyboard (“I can type on a phone, but I type 100 words a minute on a regular keyboard so that’s always gonna be my preference.”), most prefer or at least default to using an on-screen keyboard (8). One person (1) would take any kind of keyboard access “even if it’s an on-screen keyboard” but finds a physical keyboard better for their wrists and hands. Within those who prefer on-screen keyboards there was some variation in preferences of how that keyboard is implemented. Some (3) prefer a keyboard embedded within the application itself. One participant (1) said this is preferred because a native on-screen keyboard would mean “needing to swap motor plans back and forth a lot” (1) [^7] Motor planning is the process of learning how to perform an action or series of actions—or as one participant put it: “Your body knowing how to body.. A couple participants (2) had gotten around this by embedding a keyboard in the homepage of their AAC as regular buttons (for example, Figure 1). One participant (1) finds this so important that:
The biggest thing I looked for when I was initially selecting an app was the ability to add a keyboard to the homepage.
@@
[…]
A picture of an iPad with the program Proloquo2Go. The program has a big empty rectangle at the top for seeing the text that the user is currently writing to say. It is currently empty in this picture. The rest of the application is a giant grid of uniform buttons. There are categories of buttons at the very bottom of someone’s board. The home screen of their app has a lot of common words with stick figure symbols to correspond to the words. On the bottom of the grid (above the categories) there is an on-screen keyboard laid out in yellow buttons.
@@ 5.1.2. Symbol- vs text-based AAC
On the other hand, some participants have a preference for symbol-based AAC. One person (1) said that symbol-based apps work better for their brain but are much slower. Another person (1) said that they prefer having some images, although it’s not a requirement. A few participants (3) explicitly mentioned the difference they feel when using word- vs symbol-based AAC. They find typing faster (2) but compared to symbol-based buttons, they say typing creates “more mental effort” (1) or can be a “less calm” experience (1). One participant (1) will use symbol-based AAC around people because “there’s less pressure for me to find the correct words, and I have that extra support.
I like [button-based] Proloquo, though, because it forces me to slow down and think about what I’m trying to say. And for now, [typing-based] ClaroCom, when I have a thought already and need to get it out.
Some participants (7) spoke about supplementing or mixing symbol usage with typing, which can include things like using symbol-based AAC with Zoom’s chat (1) or completely mixing the two types of AAC (“Future AAC that has both symbols and typing and a vocabulary designed for autistic adults would be very empowering.”). Currently, having applications focus solely on symbol-based or text-based input slows users down.
[I dislike] symbols and typing being so completely separate in different apps. Having multiple apps is good and choice is good but sometimes its just like I don’t want to have to keep switching, and the typing in the symbol app just isn’t sufficient for what I need.
Others will actively avoid symbol-based AAC (2), although one participant (1) would accept an AAC that uses buttons without symbols. One participant said that since they see words in their head before they say them, it makes working with pictures more difficult.
5.1.3. Flexibility with vocabulary
One downside to using symbol-based AAC is that it could be limiting in the amount of vocabulary available. Seven participants (7) look for applications with high flexibility in terms of vocabulary, which may include an adult-level vocabulary (2), vocabulary for specific identity groups (1) or occupations (1), or an overall comprehensive lexicon (all 7 listed above). One participant (1) likes to use text-based AAC to see what they’re missing from their symbols. A couple of participants (2) have had experiences with older AAC with severely limited vocabulary, which they now know to avoid.
@@ 5.1.4. Layout
There were a lot of different preferences for how users want their buttons to be laid out. Some of these users wish existed in their preferred app or wish existed in general, others already exist and could be the reason why they chose the app that they use. Important aspects to consider for layouts include:
  • Grid size: The ability to set the number of buttons that appear on a single page (3) or pages that have different sized grids (1).
  • Organizing words: The order or groupings of what words appear near other words, such as alphabetically (1), semantically/associated (3), grammar-based (2), or organized manually/ color-coded (3). Two participants (2) appreciate a motor planning-based layout to help them know where words can be found. One participant (1) modified their layout to hide buttons that are used only in specific situations: “I think I keep it hidden cuz I don’t travel very much”.
  • Easy-to-find words: A layout that prevents the user from navigating folders or scrolling a lot to find a specific word (2).
  • Easy to edit: Multiple buttons can be changed or moved at the same time (2).
  • Button size: Either buttons that adapt to fill the page that they’re on (1) or buttons of different sizes on the same page (1).
  • Good core board: The app starts with a good layout that needs minimal adapting (2).
  • Consistent organization: Having the same layout across applications/devices (2).
5.1.5. Predictive text
Participants also enjoy input that aids them with drafting what they want to say, whether it be auto-complete/word suggestion (4), or helping them find the button they’re looking for via a search bar (2). Unfortunately, writing prediction is not available in symbol-based apps (1) nor when typing on a keyboard (1). A couple participants (2) have noticed that the predictive text is coming from their native keyboard rather than their app, although users who have Proloquo (1) see that it suggests related words next to the normal grid of words.
Usually these predictions are tailored to a user, which can be useful but also concerning, especially when it’s saving context-specific or misspelled words. One participant (1) suggests that future AAC could have selective forgetting of particular words.
I’ve got it programmed to save [words, and it picks up on things that are part of my] dialect like ‘wanna’ or ‘gimme’ […] I tell it to learn all those words, and it also learns anything I’m, say, telling a doctor. So, you know, it does save those words and phrases, and sometimes I misspell a word, and it’ll bring it up the wrong word to tell me […] But the ability to maybe tell it ‘forget this particular misspelling’ would be useful.
5.1.6. Pre-programed phrases
There are some features that can speed up communication. Participants (10) spoke about saving whole phrases and commonly-used messages ahead of time to be used for multiple situations.
I also like that with the phone I can pre-program phrases before I enter a situation, so if I’m gonna order, I don’t have to type it as I go, I can type it all in and hand it to them. Or I can have some emergency phrases about, you know, please call this person if I need help.
A similar feature is the quickfire reply (1), tapping on a word or phrase to immediately say the word without drafting. This can be useful for a quick yes/no response or for sentences that you know won’t change but are tiresome to type out each time, such as self introductions.
Note that quickfire replies are slightly different from speaking the words as the message is being drafted, which participants (2) disliked—one of them called it “infantilizing”.
I also hate apps that speak every word as you hit the button or speaks each sentence as you type it. It messes me up a lot and distracts me. I need to be able to turn that off so it only speaks the message when I’m finished.
Furthermore, when they do use pre-programmed messages, the person from the above quote prefers to have the option to edit it in the moment before saying it.
@@ 5.1.7. Indirect selection
The “default” input method for symbol-based AAC appears to be direct select. That is, tapping directly on the button of the word the user wishes to say. However, even in a group such autistic adults who are not necessarily using AAC due to a physical disability, some users still require or prefer an alternative mode of input. Sometimes participants will switch to an indirect select input type because a dynamic disability such as seizures (1) or hemiparesis (1) prevents them from easily using direct select.
Examples of indirect selection include:
  • Eye Gaze: Using one’s eyes to select words. Some participants are interested in or have tried using eye gaze as input: one participant (1) uses eye gaze regularly, another participant (1) would like to try it but doesn’t know how, and a third participant (1) has tried it but it gives them migraines.
  • Automatic Scanning: One button is highlighted at a time with the highlight automatically going through the buttons until the user selects one.
  • Step Scanning: A form of scanning with two switches. One button is highlighted at a time with the user pressing a switch to proceed to the next button. Another switch is hit to make a selection.
One participant (1) will use scanning or eye gaze when they have poor motor control. Their step scanning setup includes setting the “top half of the screen to select and the bottom half to advance to the next selection”. Another participant (1) wishes they could step scan with three switches instead of two: “forward, backward, and select, because I sometimes need to not be using direct select[,] and eyegaze and switch select [i.e., step scanning] both don’t work for me and it is exhausting to use a mouse”.
5.2. Output options
The two main themes that popped up in terms of output options were having the ability to show text to conversation partners and to have customizable text-to-speech (TTS)—also called speech synthesis—that fit the user’s identity. TTS & speech synthesis are the technical terms for the technology that speaks a selection of text aloud. This enables the user to share their thoughts with an entire group instead of being confined to showing a screen or piece of paper. We’ll go into these two themes (Showing text and Voice identity) in more detail in this section, but there were other interesting features that our participants mentioned either as something they have experienced and liked or would like to see in future AAC applications. These include:
  1. (1)
    A ”hold my thought” feature that saved what they were inputting so that could offer a quick reply to what was being said in the conversation and then return to what they were drafting (2).
  2. (2)
    Having the TTS begin reading out the start of a message while the remainder of it was being written (1).
  3. (3)
    Being able to quickly repeat (via TTS) what they had just said (1).
  4. (4)
    Teaching the TTS how to pronounce certain words correctly (1).
@@ 5.2.1. Showing text
Many participants (8) use the text display feature of their AAC app instead of, or in conjunction with, the TTS. This can stem from conversation participants unable to understand the synthesized voice and/or because the voice couldn’t be loud enough (6). Others prefer to show text to communicate most or all of the time (3) or switch between the different modes (1). The feature itself should also be flexible, such as with an easy adjustment of font size to show people from farther away (2) or the ability to flip the text around to show someone across from the AAC user (1). We discuss some of the limitations of showing text in Section 5.7.3.
@@ 5.2.2. Voice identity
A person’s identity is strongly tied to their voice. The availability of choice for voices (1) was an extremely important aspect for several participants, particularly not hidden behind a paywall or at least have the ability to be previewed before purchasing (1).
More customizable voices, for sure. It’s very personal and having the voice that matches every other person who uses AAC is very disempowering.
A couple participants (2) just want higher-quality voices, while others (1) know exactly what voice brands to look for (Acapela [^8] https://www.acapela-group.com/). One participant (1) has been using AAC for such a long time that they just want to use the same voice in order to sound consistent.
Although one participant (1) was happy that some voice packs were beginning to include a variety of different types of voices, one of the most common complaints we saw was that there aren’t enough nonbinary or middle-pitch voice options (4), and this becomes even more limiting when the app (such as TD Snap) limits basic customization such as adjusting the pitch (3). The only person who uses he/him pronouns out of our participants dislikes the men’s voices and prefers to pitch down child/teen TTS. Even if we were to sample their speaking voice to create a TTS voice, it might not match what they actually want to sound like:
I think the most important thing I look for is a compatible voice pack. As I mentioned before, I am trans, but I deal with age dysphoria as well, which is part of why I like [this voice]. Having a voice that sounds right is, therefore really, really important.
5.3. Selecting or adapting AAC for a good fit
Beyond input/output, participants have other criteria for selecting the AAC app that is right for them. In this category, we will go over themes around applications that fit the needs and preferences of the user.
Our participants described wanting to have a single or a few AAC application(s) that were flexible in design and would have the option to be used in multiple ways to suit their needs. This category describes themes where participants expressed the need for robust, adaptable features for their AAC so that they weren’t constrained to using an app that didn’t fit what they needed.
@@ 5.3.1. Aging up AAC
Participants (5) were of the opinion that AAC needs to be tailored for autistic adults, especially since they’re often targeted toward children.
Future AAC that has both symbols and typing and a vocabulary designed for autistic adults would be very empowering. Many AAC apps feel like they’re made for kids or students, and it feels infantilizing.
Two participants (2) described how the vocabulary available in the AAC tools had to be modified for adults. Another participant (1) felt that people who used symbol-based AAC as children might appreciate if there was a feature to switch over from images to text as they grow up. Someone (1) found that apps that report word frequency usage (“what words I need to practice or whatever”) was not a useful feature for adults who already know the language and was seen as another way of logging data. Another person (1) wished that there was more access to speech-language pathologists as an adult so that they might get more help for their AAC use.
@@ 5.3.2. Consolidation of features
Three participants (3) said that consolidating features from various apps into a single platform would make things more convenient, eliminating the need to constantly switch between different applications. Features are split across multiple applications. For instance, in one application, users could place buttons of different sizes on a single page, while another application would offer the ability to color-code items, which would be lacking in the first application. One of the participants (1) spoke about wanting to have all of the important features in one application.
There are things I like about each of the apps I use, but it would be nice to get all the features in one app. For instance, like I mentioned it’s a pain to use Proloquo2Go when you’re saving messages in advance, because you can’t easily see the whole message and everything later on. I also really wish that I could use Proloquo2Go on Swedish which I cannot. It would be nice to be able to have buttons in varying sizes on a single page, which you can do in TD Snap but not in Proloquo2Go and another feature that is missing from Proloquo2Go, which is in TD Snap is the chat feature, where if you’re in the middle of composing a message and someone interrupts that with another question, you can hit the chat button, compose a reply to the second thing, and then go back to composing the first message which had been saved where you left it.
There are also phone features that participants wished could be better integrated into their AAC applications to make their lives easier. These features included texting (1), phone calls (2), and the ability to easier play AAC audio from the phone speaker when headphones are being used (1).
@@ 5.3.3. General customization
All of our participants (12) expressed various levels of interest in overall customization and ease of use.
I like being able to adjust pretty much every feature really. Whether that’s the appearance of the buttons, how the grammar support behaves, how to navigate into and out of folders and so on.
Participants also expressed various preferences for how their ideal AAC application should look:
  • Preferred symbol library (4) - Boardmaker symbols, SymbolStix symbols (2), one with multiple skin tones (1), or having multiple library options—one participant (1) considers this when recommending an app to others.
  • Not overwhelming (7) - Interface that is not visually overwhelming (2), such as a clear layout and organization with fewer possibilities (4), or no bright colors (3)—especially when they are already feeling overwhelmed.
  • Easy-to-see symbols (2) - Symbols are high contrast (1). Another participant (1) avoids symbol-based AAC as someone who is low-vision.
  • Other visual customization (2) - Such as font selection (1) or dark mode for people with light sensitivity (1).
5.4. When to Start, Swap, or Stop AAC
Researchers still don’t know why autistic people may have trouble speaking orally, although some suggest that it could be anxiety-based (Muris and Ollendick, 2021; Ludlow et al., 2023)—although it’s hard to say whether anxiety is creating the difficulties or that having difficulties speaking orally gives people anxiety.
Melfsen et al. (2021)’s “unsafe world model” of selective mutism extends beyond anxiety and describes the lack of oral speech as a “shut down” of the nervous system from a perceived lack of safety. In this model, over-stimulation (among other things) can cause the perception of being unsafe. Related to this, two recurrent themes that came up in our study were the aspects of Trust of other people and the use of AAC to help them with Accommodating to current needs, which we will discuss in this section.
A third theory is that selective mutism could come from a potential weakness in neural auditory pathways (Muchnik et al., 2013; Henkin and Bar-Haim, 2015). In this section we will also discuss reasons why people might switch from one type of AAC to another—outside of input selection (§5.1) or output selection (§5.2).

### v3

@@ 25-34 Semi-speaking 5-9 Proloquo4Text,
25-34 Mostly non-speaking $\leq$ 1 Proloquo,
@@ 35-44 Unreliable speech $\leq$ 1 TD Snap,
[…]
Table 1 is a list of participants and their varying AAC use. Most AAC applications listed are in red (denoting applications), with the exception of writing, ASL, UbiDuo, Relay, texting/messaging, email, and many low-tech options. Some participants have AACs that are underlined (denoting what they use the most), but others don’t have a most-commonly-used AAC.
@@ 5. Results
This section delves into the categories identified during the qualitative analysis of the transcripts. We use a numerical value enclosed in parentheses to indicate the number of participants supporting that particular theme (e.g., ”(3)” signifies ”three interviewees”). We will not be using any individual IDs due to the close-knit community of AAC users. Because of this, we will refer to everyone by ”they/them”, although we collected their pronouns during the interview. Direct quotes will be surrounded by quotation marks and italicized. The complete list of all 12 participants can be found in Table 1.
We have identified 9 main categories of themes. Because our participants had a lot to say about what they look for in an AAC application (RQ2), we split this into three categories: Input Flexibility5.1), Output Flexibility5.2), and Selecting or Adapting AAC5.3). Benefits (RQ1) can be found in Section 5.5. Challenges (RQ1) can be found in Stumbling blocks for continued use5.7). For RQ4, we have the category: Control of Communication5.8). Themes surrounding our participants’ identities (RQ3) can be found across categories. Additionally, two categories emerged from the interviews: Contexts for AAC Use5.4) and Access as an Adult5.6).
5.1. Input Flexibility
All 12 participants had preferences of what input type they preferred or would like to see in future AAC applications. In this section, we discuss the use of keyboards (§5.1.1), the differences felt between symbol-based and text-based AAC (§5.1.2), and aspects that make input easier such as vocabulary (§5.1.3) & layout flexibility (§5.1.4), predictive text/pre-programmed phrases (§5.1.5), and button selection interfaces (§5.1.6).
@@ 5.1.1. Typing
Since our participants were recruited because they were autistic, they all had varying degrees of motor control, which often changes depending on their physical, emotional, or mental state. That said, all participants (12) wanted some type of typing input option to use at least occasionally. This could range from full-time use to simply augmenting their vocabulary for when they don’t have an appropriate button. Although a few (3) prefer a physical keyboard (“I can type on a phone, but I type 100 words a minute on a regular keyboard so that’s always gonna be my preference.”), most prefer or at least default to using an on-screen keyboard (8). One person (1) would take any kind of keyboard access “even if it’s an on-screen keyboard” but finds a physical keyboard better for their wrists and hands. Within those who prefer on-screen keyboards there was some variation in preferences of how that keyboard is implemented. Some (3) prefer a keyboard embedded within the application itself. One participant (1) said this is preferred because an OS-native on-screen keyboard would mean “needing to swap motor plans back and forth a lot” (1) [^8] Motor planning is the process of learning how to perform an action or series of actions—or as one participant put it: “Your body knowing how to body.. A couple participants (2) had gotten around this by embedding a keyboard in the homepage of their AAC as regular buttons (for example, Figure 2).
Figure 2 is a picture of an iPad with the program Proloquo2Go. The program shows a grid of buttons with varying symbols denoting words. On the bottom of the grid there is a QWERTY keyboard laid out in yellow buttons (to visually pop compared to the usually-white button backgrounds).
@@ 5.1.2. Symbol- vs text-based AAC
On the other hand, some participants have a preference for symbol-based AAC. One person (1) said that symbol-based apps work better for their brain but are much slower. Another person (1) said that they prefer having some images, although it is not a requirement. A few participants (4) explicitly mentioned the difference they feel when using word- vs symbol-based AAC. They find typing to be faster (2) but more taxing (2) — creating “more mental effort” or being “less calm”. Meanwhile, regarding symbol-based AAC, one participant said “there’s less pressure for me to find the correct words,” while another said “it forces me to slow down and think about what I’m trying to say.” Others will actively avoid symbol-based AAC (2), or allow buttons as long as they don’t have symbols (1). Holyfield (2021) has shown that even pre-literate autistic children take about the same amount of time to learn symbol-based or text-based AAC, so perhaps the choice is personal preference.
Some participants (7) suggested supplementing or mixing symbol usage with typing, which can include things like using symbol-based AAC with Zoom’s chat (1) or completely mixing the two types of AAC (“Future AAC that has both symbols and typing and a vocabulary designed for autistic adults would be very empowering.”). Having applications focus solely on symbol-based or text-based input creates unnecessary friction:
[I dislike] symbols and typing being so completely separate in different apps. Having multiple apps is good and choice is good but […] I don’t want to have to keep switching, and the typing in the symbol app just isn’t sufficient for what I need.
5.1.3. Vocabulary
One downside to using symbol-based AAC is that it could be limiting in the amount of vocabulary available. As adults who have been using mouthwords for years, our participants have very large vocabularies. A couple of participants (2) have had experiences with older AAC with severely limited vocabulary. Seven of our participants (7) look for applications with high flexibility in terms of vocabulary, which may include an adult-level vocabulary (2), vocabulary for specific groups of people (2), or an overall comprehensive lexicon (7). The integration of adult or job-specific vocabulary has also been suggested by McNaughton and Bryen (2007).
@@ 5.1.4. Layout
There were a lot of different preferences for how users want their buttons to be laid out. Some of these our participants wish existed in their preferred app or wish existed in general, others already exist and could be the reason why they chose the app that they use. Our participants (7) had different opinions regarding aspects of the app’s grid size (4), word organization (5), ease of use (4), button size (2), quality of the core board (2), and ability to use the same layout across applications/devices (2).
5.1.5. Predictive & pre-programmed text
Participants also enjoy input that aids them with writing what they want to say, whether it be auto-complete/word suggestion (4), or helping them find the button they are looking for via a search bar (2). However, our participants echoed some of the issues of existing AAC prediction, such as the system unhelpfully storing misspellings (1) (McKillop, 2018) or suggesting words that were inappropriate for the social context (1) (McKillop, 2018; Valencia et al., 2023). Thankfully, next-word prediction for AAC is frequently researched (Schadle, 2004; Trnka et al., 2007; Wiegand and Patel, 2012; Freedman et al., 2013; Adhikary et al., 2019) – as well as contextual or discourse-based prediction (Freedman et al., 2013; Mitchell and Sproat, 2012; Perkoff, 2021), personalized language (Freedman et al., 2013; Di Mascio et al., 2019; Li et al., 2022; Valencia et al., 2023), or generating sentences from a set of keywords (Cai et al., 2022; Kumar et al., 2022). More work is needed for determining how predictive text is displayed (Yusufali et al., 2023b) and how to guarantee that the predictive text does not undermine autonomy (Valencia et al., 2023).
Many participants (10) also used features to pre-program whole phrases and commonly-used messages ahead of time.
@@ 5.1.6. Indirect selection
The “default” input method for symbol-based AAC appears to be direct select. That is, tapping directly on the button of the word the user wishes to say. However, even in a group such autistic adults who are not necessarily using AAC due to a physical disability, some users still require or prefer an alternative mode of input. Sometimes participants (3) will switch to an indirect select input type (e.g., eye gaze, automatic scanning, or step scanning) because a dynamic disability such as seizures (1) or hemiparesis (1) prevents them from easily using direct select.
5.2. Output Flexibility
The two main themes that popped up in terms of desired output options were having the ability to show text to conversation partners and to have customizable text-to-speech (TTS)—also called speech synthesis—that fits the user’s identity. TTS enables the user to share their thoughts with an entire group instead of being confined to showing a screen or piece of paper. We will go into these two themes (Showing text and Voice identity) in more detail in this section. (We will talk about issues with showing text and speech synthesis in Sections 5.7.3 & 5.7.4, respectively.)
@@ 5.2.1. Showing text
Many participants (8) use the text display feature of their AAC app instead of, or in conjunction with, the TTS. This can stem from conversation participants unable to understand the synthesized voice and/or because the voice couldn’t be loud enough (6). Others prefer to communicate by showing text most or all of the time (3) or switch between speech and text (1).
@@ 5.2.2. Voice identity
A person’s identity is strongly tied to their voice (Nathanson, 2017), and having a unique, personalized AAC voice can be empowering (Engelke, 2013; Mills et al., 2014; McGettigan, 2015; Patel and Threats, 2016; Zeffate and Elhari, 2023; Preece et al., 2024). Our participants expressed an interest in having access to options for different types of voices (2) or access to specific voices from voice packs that they have used in the past (2).
On top of this, many of our participants were transgender or gender nonconforming and wanted voices to reflect this. Although one participant was happy some voice packs were beginning to include a variety of voices, a common complaint that we saw was that there are not enough nonbinary or middle-pitch voice options (4), and this becomes even more limiting when the app prohibits even basic customizations such as adjusting the pitch (3). The only person who uses he/him pronouns out of our participants dislikes the men’s voices and prefers to pitch down child/teen TTS. Even if we were to sample their speaking voice to create a TTS voice, it might not match what they actually want to sound like, making voices based on the user a less desirable option: “ I am trans, but I deal with age dysphoria as well, which is part of why I like [this voice]. Having a voice that sounds right is, therefore really, really important.
5.3. Selecting or Adapting AAC
Beyond input/output, participants have other criteria for selecting the AAC app that is right for them. In this category, we will go over themes around applications that fit the needs and preferences of the user. Our participants described wanting to have a single or a few AAC application(s) that were flexible in design and would have the option to be used in multiple ways to suit their needs. This category describes themes where participants expressed the need for robust, adaptable features for their AAC so that they weren’t constrained to using an app that didn’t fit what they needed. These include: Aging up AAC5.3.1), Consolidation of features5.3.2), and General customization5.3.3).
@@ 5.3.1. Aging up AAC
Many AAC apps feel like they’re made for kids or students, and it feels infantilizing.
Participants (5) were of the opinion that AAC needs to be tailored for autistic adults, especially since they are often targeted toward children. Four participants (4) had suggestions for how AAC could be modified for adults: providing the user access to more adult vocabulary (2; also mentioned in §5.1.3), which McNaughton and Bryen (2007) have also found; switching from symbol-based to text-based AAC as someone grows up, or combining the two modalities (1); and no recording of word frequency usage [^9] Word usage is usually sent to a speech-language pathologist when learning how to use AAC to inform them “what words I need to practice or whatever”., especially since this can be seen as another way of logging data (1). Further work should be done to study how the autistic adult community believes AAC can “age up”.
@@ 5.3.2. Consolidation of features
Similar to other AAC users (Obiorah et al., 2021; Dai et al., 2022), our participants wish AAC could fit into their existing tech ecosystem. Three participants (3) said that consolidating features from various apps into a single platform would make things more convenient, eliminating the need to constantly switch between different applications. There are also existing phone features, such as texting or voice calling, that participants wished could be better integrated into their AAC applications to make users’ lives easier (3). Other participants had a dedicated device for AAC (5) or expressed interest in getting one (1).
@@ 5.3.3. General customization
All of our participants (12) expressed various levels of interest in overall customization and ease of use. This makes sense considering autistic people have different types of needs. One of the most commonly-expressed preferences was that participants wanted their ideal AAC application to not feel overwhelming (7). Depending on the person, this could mean an interface that is not visually overwhelming (2), such as a clear layout and organization with fewer possibilities (4), or no bright colors (3)—especially when they are already feeling overwhelmed.
5.4. Contexts for AAC Use
There are a few theories about why autistic people may have trouble speaking orally. Some suggest that it could be anxiety-based (Muris and Ollendick, 2021; Ludlow et al., 2023)—although it is hard to say whether anxiety is creating the difficulties or that having difficulties speaking orally gives people anxiety. Melfsen et al. (2021)’s “unsafe world model” of selective mutism extends beyond anxiety and describes the lack of oral speech as a “shut down” of the nervous system from a perceived lack of safety. In this model, over-stimulation (among other things) can cause the perception of being unsafe. Related to this, two recurrent themes that came up in our study were the aspects of Trust of other people and the use of AAC to help them with Accommodating to current needs, which we will discuss in this section.